OpenStudy (anonymous):

Graph the two equations to solve the system. What is the ordered pair solution? x + y = 6 and x - y = 4

OpenStudy (anonymous):

1 x + 1 y = 6 .............1 1 x -1 y = 4 .............2 equation1 when x= 0 y= 6 ( 0 , 6 ) when y = 0 x= 6 ( 6 , 0 ) Equation 2 when x= 0 y= -4 ( 0 , -4 )

OpenStudy (anonymous):

when y = 0 x= 4 ( 4 , 0 )

OpenStudy (anonymous):

The answer choices are - (4,2) (6,2) No Solution (5,1)

OpenStudy (anonymous):

the answer is (2,4)